Magnesium sulfate is often used in first-aid hot packs, giving off heat when dissolved in water. A coffee-cup calorimeter at 25 degrees celsius contains 15.00 mL of water at 25 degrees celsius. A 2.00-g sample of magnesium sulfate is dissolved in the water and 1.51 kJ of heat are evolved. What is the final temperature (celsius) of the solution when density of water is 1.00 g/mL and the specific heat of water is 4.18 J/g-degree celsius?
1
Expert's answer
2020-09-15T13:50:39-0400

QUESTION # 133243

Magnesium sulfate is often used in first-aid hot packs, giving off heat when dissolved in water. A coffee-cup calorimeter at 25 degrees celsius contains 15.00 mL of water at 25 degrees celsius. A 2.00-g sample of magnesium sulfate is dissolved in the water and 1.51 kJ of heat are evolved. What is the final temperature (celsius) of the solution when density of water is 1.00 g/mL and the specific heat of water is 4.18 J/g-degree celsius?

ANSWER

ΔHsolution  = MCΔT

M = 15mL×1.00g/mL =15g

C = 4.18 J/g-degree Celsius

ΔT= Change in temperature = final temperature -initial temperature

ΔHsolution =1.51kJ=1.51×1000=1510J

Given, ΔHsolution  = MCΔT

             1510J = 15g×4.18 J/g-℃×ΔT

              1510J = 62.7J/℃×ΔT

               ΔT= 1510J÷62.7J/℃

                   = 24.08℃

ΔT= Change in temperature = final temperature -initial temperature

24.08℃ = final temperature-25 ℃

final temperature=24.08℃+25 ℃

                           =49.08℃

                           =49℃